OpenStudy (anonymous):

double check my work?? If y = -7 when x = -3, find y when x = 9. I got y=21

zepdrix (zepdrix):

ya sounds good c:

zepdrix (zepdrix):

y is some multiple of x... so y=kx. -7=k(-3) --> k=7/3 y=(7/3)(9)=21